A Client has $160,000.00 getting 3% interest. Client would like to Withdrawal $600 per month. How long would the $160K last at that Withdrawal/Interest Rate?
Thank you!

Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

Is the 3% compounded monthly? Quarterly? Annually? I will assume it's compounded monthly
:
n = number of withdrawals(we will find this number)
A = present value 160,000
P = monthly payment 600
i = interest monthly 0.03/12 = 0.0025
:
n = log[P/(P-A*i)]/log(1+i)
n = log[600/(200-15000*.0025)]/log(1.0025) =
n = log(600/162.5)/0.0025
n = 0.5673/0.0025 = 226.92
:
I think the numbers are correct. Good luck.

Answer by MathTherapy(10552)

A Client has $160,000.00 getting 3% interest. Client would like to Withdrawal $600 per month. How long would the $160K last at that Withdrawal/Interest Rate?
Thank you!